OpenStudy (anonymous):

Is there a function that resembles the \(\text{sqrt}\) function but whose limit as \(x\to\infty\) exists?

OpenStudy (turingtest):

not sure what you mean by "resembles" the sqrt function but\[\large\lim_{n\to\infty}{n\over\sqrt[n]{n!}}=e\]

OpenStudy (anonymous):

I am looking for something like this.

OpenStudy (anonymous):

A function like the red one. That is, a function that converges as \(x\to\infty\), but whose inverse diverges as \(x\to\infty\).

OpenStudy (turingtest):

what about \(y=e^{-x}\)? doesn't that fulfill the requirement above?

OpenStudy (anonymous):

That one will do. Thank you very much! :)

OpenStudy (turingtest):

yay, I answered one of your questions finally ;)